$7,494 $2,633.75 $5,925.93 $10,000 4. New equipment used in construction with a cost basis of $60,000 with a zero salvage value at the end of its useful life of 6 years. Use the Double Declining method to calculate the depreciation charge in year 4.
Answer #1

COST OF EQUIPMENT $ 60,000

ZERO SALVAGE

USEFUL LIFE= 6 YEARS

RATE OF DEPRECIATION IS TAKEN AS DOUBLE OF STRAIGHT LINE METHOD.

UNDER STRAIGHT LINE METHOD DEPRECIATION WOULD BE

COST OF EQUIPMENT/USEFUL LIFE 60000/6 = 10000

10000/60000= 0.166

0.166*100 = 16.66%

WHEN WE DOUBLE IT WE GET 33.3%

SO 33.3% IS OUR RATE OF INTEREST

YEAR BEGINNING BOOK VALUE DEP % DEP AMOUNT TOTAL DEP AMOUNT ENDING BOOK VALUE

1. $60,000 33.3% $20,000 $20,000 $40,000

2. $40,000 33.3% $13,333 $33,333 $26,667

3. $26,667 33.3% $8,889 $42,222 $17,778

4. $17,778 33.3% $5,926 $48,148 $11,852

5. $11,852 33.3% $3,951 $ 52,099 $7,901

6. $ 7,901 100% $7901 $60,000 00

DEPRECIATION IS CALCULATED EACH YEAR AT THE BEGINNING BOOK VALUE

TOTAL DEPRECIATION AMOUNT IS THE ACCUMULATED DEPRECIATION OVER THE YEARS

ENDING BOOK VALUE IS AMOUNT OF DEPRECIATION LESS FROM THE BOOK VALUE OF THAT YEAR,WHICH BECOMES BEGINNING VALUE OF THE NEXT YEAR

AS PER ABOVE CALCULATIONS,THE VALUE OF THE FOURTH YEAR IS $5,926

CORRECT ANSWERS FROM GIVEN OPTIONS IS $5,925.93